# Improving Efficiency of Sampling-based Motion Planning via Message-Passing Monte Carlo

### v1

@@ V-B OMPL and UR5 benchmarks
Cubicle SE(3) Uniform 00 $56.4\pm 5.0$ 00 $112.9\pm 7.2$ $2$ $221.8\pm 13.6$ $26$ $412.6\pm 60.8$
Halton 00 $57.4\pm 3.4$ 00 $112.3\pm 4.8$ $4$ $221.1\pm 8.9$ $36$ $382.4\pm 98.9$
MPMC 00 $58.8\pm 3.3$ 00 $112.3\pm 3.6$ $10$ $221.2\pm 14.5$ $46$ $367.4\pm 100.5$
Twistycool SE(3) Uniform $8$ $89.2\pm 14.7$ $4$ $180.4\pm 22.3$ $26$ $332.7\pm 71.6$ $70$ $485.0\pm 221.7$
Halton $6$ $89.9\pm 10.8$ $16$ $171.4\pm 30.2$ $20$ $336.0\pm 63.0$ $56$ $584.1\pm 193.4$
MPMC $10$ $89.5\pm 11.5$ $8$ $175.2\pm 27.6$ $30$ $327.7\pm 67.1$ $72$ $499.3\pm 212.7$
@@
MPMC $48$ $15.4\pm 12.4$ $60$ $31.6\pm 24.4$ $68$ $69.3\pm 45.5$ $84$ $161.0\pm 71.0$
@@
Halton $20$ $5.3\pm 6.5$ $32$ $15.0\pm 14.2$ $34$ $27.6\pm 27.3$ $28$ $42.4\pm 52.3$
MPMC $56$ $10.3\pm 6.3$ $64$ $22.5\pm 12.7$ $76$ $40.6\pm 22.6$ $96$ $103.1\pm 20.1$
Hypercube $\mathbb{R}^{10}$ Uniform $2$ $7.0\pm 2.0$ $12$ $9.7\pm 2.5$ $62$ $17.7\pm 4.3$ $88$ $33.9\pm 5.5$
Halton $2$ $6.6\pm 1.8$ $12$ $10.4\pm 2.3$ $56$ $17.7\pm 2.9$ $98$ $32.6\pm 3.4$
MPMC 00 $6.3\pm 1.7$ $16$ $10.7\pm 2.4$ $80$ $18.9\pm 1.9$ $100$ $31.1\pm 2.9$
Kinematic Chain $\mathbb{R}^{10}$ Uniform $4$ $30.2\pm 4.8$ $12$ $57.5\pm 9.6$ $22$ $103.5\pm 26.3$ $54$ $172.5\pm 64.9$
Halton $2$ $30.0\pm 4.7$ $14$ $55.5\pm 9.1$ $22$ $103.7\pm 21.4$ $62$ $156.0\pm 64.5$
MPMC $6$ $31.5\pm 4.3$ $24$ $51.8\pm 7.5$ $26$ $102.3\pm 21.5$ $60$ $161.7\pm 71.9$
UR5 $\mathbb{R}^{5}$ Uniform 00 $47.6\pm 5.8$ $10$ $92.2\pm 7.8$ $6$ $184.3\pm 11.5$ $12$ $368.6\pm 16.1$
Halton $2$ $49.4\pm 6.9$ $4$ $93.1\pm 7.5$ $14$ $185.8\pm 10.8$ $18$ $369.2\pm 15.6$
MPMC $6$ $49.1\pm 3.0$ $31$ $97.4\pm 3.5$ $75$ $181.4\pm 3.8$ $81$ $360.6\pm 4.3$
VI DISCUSSION
This work introduces the application of MPMC neural network point set generation in motion planning, offering a novel, unbiased approach to sampling. Our proposed method includes a custom training objective specifically designed for high-dimensional configuration spaces, ensuring that the generated points are optimized for complex planning tasks. Also, we provide theoretical proof that the MPMC point sets offer a tighter upper bound on the distance from the optimal path.
@@ VI DISCUSSION
One limitation of our approach is the need to retrain the MPMC model for each specific number of points $N$ and dimensionality $d$ to ensure optimal performance. This retraining requirement can be computationally intensive, particularly for applications where the configuration space frequently changes. Future work will focus on refining the neural network architecture and training procedure to achieve generalization across both $N$ and $d$, reducing the need for retraining and making the method more flexible across different planning scenarios.
@@
Moreover, while we have demonstrated the benefits of MPMC within the PRM framework, future research will explore the potential advantages of integrating this technique into more sophisticated sampling-based planners, potentially extending its impact across a broader class of planning algorithms. Finally, an exciting avenue for future exploration lies in adapting our technique to compute-critical applications, such as acrobatic flight or high-speed driving, where planners like the Model Predictive Path Integral (MPPI) [41] require real-time performance and could greatly benefit from efficient, high-quality sampling methods.

### v2

@@ 5.2 OMPL and UR5 benchmarks
Cubicle SE(3) Uniform 0 $56.4\pm 5.0$ 0 $112.9\pm 7.2$ 22 $221.8\pm 13.6$ $26$ $412.6\pm 60.8$
Halton 0 $57.4\pm 3.4$ 0 $112.3\pm 4.8$ 44 $221.1\pm 8.9$ $36$ $382.4\pm 98.9$
MPMC 0 $58.8\pm 3.3$ 0 $112.3\pm 3.6$ $\mathbf{10}$ $221.2\pm 14.5$ $\mathbf{46}$ $367.4\pm 100.5$
Twistycool SE(3) Uniform 88 $89.2\pm 14.7$ 44 $180.4\pm 22.3$ $26$ $332.7\pm 71.6$ $70$ $485.0\pm 221.7$
Halton 66 $89.9\pm 10.8$ $\mathbf{16}$ $171.4\pm 30.2$ 2020 $336.0\pm 63.0$ $56$ $584.1\pm 193.4$
MPMC $\mathbf{10}$ $9.5\pm 11.5 88$ $175.2\pm 27.6$ $\mathbf{30}$ $327.7\pm 67.1$ $\mathbf{72}$ $499.3\pm 212.7$
@@
MPMC $\mathbf{48}$ $15.4\pm 12.4$ $\mathbf{60}$ $31.6\pm 24.4$ $\mathbf{68}$ $69.3\pm 45.5$ $\mathbf{84}$ $161.0\pm 71.0$
@@
Halton 2020 $5.3\pm 6.5$ $32$ $15.0\pm 14.2$ $34$ $27.6\pm 27.3$ $28$ $42.4\pm 52.3$
MPMC $\mathbf{56}$ $10.3\pm 6.3$ $\mathbf{64}$ $22.5\pm 12.7$ $\mathbf{76}$ $40.6\pm 22.6$ $\mathbf{96}$ $103.1\pm 20.1$
Hypercube $\mathbb{R}^{10}$ Uniform $\mathbf{2}$ $7.0\pm 2.0$ $12$ $9.7\pm 2.5$ $62$ $17.7\pm 4.3$ $88$ $33.9\pm 5.5$
Halton $\mathbf{2}$ $6.6\pm 1.8$ $12$ $10.4\pm 2.3$ $56$ $17.7\pm 2.9$ $98$ $32.6\pm 3.4$
MPMC 0 $6.3\pm 1.7$ $\mathbf{16}$ $10.7\pm 2.4$ $\mathbf{80}$ $18.9\pm 1.9$ $\mathbf{100}$ $31.1\pm 2.9$
Kinematic Chain $\mathbb{R}^{10}$ Uniform 44 $30.2\pm 4.8$ $12$ $57.5\pm 9.6$ $22$ $103.5\pm 26.3$ $54$ $172.5\pm 64.9$
Halton 22 $30.0\pm 4.7$ $14$ $55.5\pm 9.1$ $22$ $103.7\pm 21.4$ $\mathbf{62}$ $156.0\pm 64.5$
MPMC $\mathbf{6}$ $31.5\pm 4.3$ $\mathbf{24}$ $51.8\pm 7.5$ $\mathbf{26}$ $102.3\pm 21.5$ $60$ $161.7\pm 71.9$
UR5 $\mathbb{R}^{5}$ Uniform 0 $47.6\pm 5.8$ $10$ $92.2\pm 7.8$ 66 $184.3\pm 11.5$ $12$ $368.6\pm 16.1$
Halton 22 $9.4\pm 6.9 44$ $93.1\pm 7.5$ $14$ $5.8\pm 10.8 1818$ $369.2\pm 15.6$
MPMC $\mathbf{6}$ $49.1\pm 3.0$ $\mathbf{31}$ $97.4\pm 3.5$ $\mathbf{75}$ $1.4\pm 3.8 𝟖𝟏\mathbf{81}$ $360.6\pm 4.3$
6 Discussion
This work introduces the application of MPMC neural network point set generation in motion planning, offering a novel, unbiased approach to sampling. Our proposed method includes a custom training objective specifically designed for high-dimensional configuration spaces, ensuring that the generated points are optimized for complex planning tasks. Moreover, we rigorously show that MPMC point sets offer a tighter upper bound on the distance from the optimal path.
@@ 6 Discussion
One limitation of our approach is the need to retrain the MPMC model for each specific number of points $N$ and dimensionality $d$ to ensure optimal performance. This retraining requirement can be computationally intensive, particularly for applications where the configuration space frequently changes. Future work will focus on refining the neural network architecture and training procedure to achieve generalization across both $N$ and $d$, reducing the need for retraining and making the method more flexible across different planning scenarios. Possible avenues to mitigate this limitation are discussed in Appendix A.2.
@@
Moreover, while we have demonstrated the benefits of MPMC within the PRM framework, future research will explore the potential advantages of integrating this technique into more sophisticated sampling-based planners, potentially extending its impact across a broader class of planning algorithms. Finally, an exciting avenue for future exploration lies in adapting our technique to compute-critical applications, such as acrobatic flight or high-speed driving, where planners require real-time performance and could greatly benefit from efficient, high-quality sampling methods.
Acknowledgments
This work was supported in part by the Postdoc.Mobility grant P500PT-217915 from the Swiss National Science Foundation, the National Science Foundation EFRI program under grant number 1830901, the Gwangju Institute of Science and Technology, and the Department of the Air Force Artificial Intelligence Accelerator and was accomplished under Cooperative Agreement Number FA8750-19-2-1000. The views and conclusions contained in this document are those of the authors and should not be interpreted as representing the official policies, either expressed or implied, of the Department of the Air Force or the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for Government purposes notwithstanding any copyright notation herein.
Appendix A Appendix
A.1 2D point set visualizations
Figure 6 depicts point sets generated by the various sampling schemes used in the 2D-maze experiments namely, Uniform, MPMC, Halton, and Sobol for different numbers of points. Uniformly sampled points exhibit large gap regions for all numbers of points. The sequential constructions of Halton and Sobol are more evenly spread but are surpassed in discrepancy terms by the learned MPMC points.
Figure 6: Points sampled on the unit square for $N\in\{32,64,128,256\}$ and sampling distributions among Uniform, Halton, Sobol, and MPMC.
A.2 MPMC point (re)generation
The solution as presented in this work requires retraining a GNN point cloud for each $(N,d)$ pair. Indeed, we focus on optimized point sets (not sequences) to maintain the clearest link between low-discrepancy sampling and improved PRM planning efficiency. There are, however, many practical solutions for mitigating this practicality obstacle and adapting this approach to sequential point generation:
  1. 1.
    For sequences over the number of points $N$, we propose training a large MPMC point set that can be transformed into a low-discrepancy sequence using a greedy approach, i.e., successively adding points that minimize the discrepancy over all choices in the point set. This further obviates the need for pruning.
  2. 2.
    Inductive graph learning is another solution to $N$ generalization, where a GNN can be trained to perform across a varying range of number of nodes (i.e., points). This approach has proven its efficacy in many other contexts.
  3. 3.
    To deal with varying $d$, a high dimensional point cloud can be used for lower dimensional problems via canonical projection of the MPMC points. Indeed, it has been shown in [11] that MPMC points yield low-discrepancy in particular for lower dimensional projections of the point sets.
A.3 Dispersion and grids
Sukharev grids are structured sampling patterns that achieve uniform coverage by placing one point at the center of each cell in an evenly spaced Cartesian grid. In $d$ dimensions, the domain $0,1]^{d}$ is divided into $k$ equal intervals along each axis, producing $k^{d}$ hypercubes, and one sample is placed at the center of each hypercube. Sukharev grids are $\mathcal{D}_{\infty}$-dispersion optimal when $N=k^{d}$, with $k$ an integer. For arbitrary $N$, however, they can be wasteful: for example, with $N=100$ and $d=3$, the best the grid can do is $4\times 4\times 4=64$ points on the lattice, leaving $36$ points that do not contribute to an improvement in dispersion. This inefficiency becomes increasingly severe in higher dimensions.
A.4 The case of non‑Euclidean manifolds
MPMC can be extended to general manifolds by developing appropriate objective functions. As an example, MPMC can be extended to $S^{n}$ by minimizing the total euclidean distance between all pairs of points. However, a general framework of MPMC on manifolds will require a separate investigation to measure the best approach and performance gains in such cases.